# AutoAct: Automatic Agent Learning from Scratch via Self-Planning

### v2

@@ 4 Results
Under identical settings, multi-agent architectures generally exhibit better performance than single-agent (ReAct vs. BOLAA, FireAct vs. AutoAct), which aligns with Simon’s theory of bounded rationality. Seemingly contrary to expectations, despite being a single-agent architecture, Chameleon outperforms BOLAA (even FireAct on ScienceQA). However, we analyze that this can be attributed to the way it leverages tools. In Chameleon, the process of deciding tool parameters is considered a form of tool invocation, and specialized few-shot prompts are designed to guide the model through this process. From this aspect, Chameleon, despite being nominally a single-agent architecture, exhibits characteristics that resemble a multi-agent architecture, which does not contradict our initial conclusion. Indeed, we can also explain from the perspective of optimizing objectives. Another well-known economic principle, Goodhart’s Law Goodhart (1984), states that “When a measure becomes a target, it ceases to be a good measure”. This implies that optimizing one objective on the same agent will inevitably harm other optimization objectives to some extent. Therefore, it is not optimal to optimize all objectives on a single agent, and a multi-agent architecture happens to address this issue. However, we analyze in §5 that excessive fine-grained division-of-labor is not the best approach and a moderate division of labor benefits group performance.
@@
Table 3 presents the performance of AutoAct on the 70b model after removing certain key processes. It can be observed that the least impactful removal is the - reflect. We investigate that in the zero-shot scenario, the model tends to be over-confident in its answers. It typically only recognizes its errors when there are obvious formatting mistakes or significant repetitions in the planning process. Consistent with previous findings, the removal of the - multi agents leads to a noticeable decrease in performance. A more exciting discovery is that the results of - multi are comparable to those of FireAct. This indirectly suggests that the trajectory quality generated by the 70b model may be no worse than that of GPT-4. As expected, the performance deteriorates after - fine-tuning, which once again confirms the previous conclusion. To demonstrate the necessity of filtering out planning error data, we specifically remove the filtering process (- filtering) to examine the performance of AutoAct. The results indicate that the damage caused by training on unfiltered data is even greater than that of - fine-tuning.
Figure 5: Case study. AutoAct (b) successfully addresses the failure in ReAct (a) by employing a more scientific combination of tools and making more accurate tool invocations. With more planning rounds, AutoAct (c) can validate its inner answers by continuing more rounds of self-verification. While this can also lead to a longer context, gradually deviating AutoAct (d) from the original question.
Figure 6: Human evaluation of trajectories generated by Llama-2-70b-chat on HotpotQA. We compare the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and the overall coherence of each trajectory.
@@ 5 Analysis
We evaluate the influence of different training data scales on the performance of self-planning in Figure 3(a-c). It can be observed that the overall performance of different models steadily improves as the training data scale increases. However, it goes to stability with minimal fluctuations once the data scale exceeds 200. We speculate that this may be due to the limited ability of naive self-instruct to boost internal knowledge of the language model. The process from self-instruct to self-planning can be seen as the distillation and rumination of internal model knowledge. As the training data increases, the knowledge which can be extracted through self-instruct decreases. Despite our efforts to filter out duplicate data, similar data will inevitably reoccur. So the mindless increase in data scale can lead to a significant surge in similar data, which undermines the benefits of increasing the data scale and ultimately makes it challenging to improve model performance or even leads to over-fitting (see the slight decrease in overall performance on 70b at 250 and 300). To further confirm the role of training data, we decouple the models from the training data and evaluate their training results on trajectories synthesized by other stronger models. From Figure 3(d-f), we can see consistent conclusions with previous findings. The performance improvement becomes increasingly challenging beyond a dataset size of 200, regardless of the size matching between the backbone model and the data-synthetic model. Therefore, maximizing the diversity of the synthesized data in the database may be a key improvement direction for AutoAct. Some previous works Xu et al. (2023); Yu et al. (2023); Li et al. (2023) have attempted to improve upon the naive self-instruct, but none of them have focused on better mobilizing the language model’s internal knowledge without external information, and we leave this for our future work.
Moderate division of labor benefits group planning performance.
@@
To explore the impact of the different granularity of self-differentiation and group planning, we further subdivide the tool agent, assigning dedicated agents to manipulate each specific tool. We contrast the performance of One agent, Three agents (vanilla AutoAct), and the Tool-Specified setting in Figure 4. It can be observed that finer task allocation does not necessarily lead to better performance. This is consistent with the findings in Qiao et al. (2023a) which indicate that multi-tool joint learning often outperforms single-tool individual learning. Therefore, appropriate differentiation (Three) can alleviate the pressure on individual agents, aligning with Simon’s principle of bounded rationality. However, excessive differentiation (Tool-Specified) not only fails to achieve better results but can sometimes even be less effective than not differentiating (One) at all. Moreover, it appears that the performance loss of tool-specific agents compared to the three-agent approach is more significant on harder problems. This is because challenging problems typically require more planning steps and higher levels of collaboration among tools. By unifying tool invocations under one agent, it becomes possible to effectively learn the interplay and interconnectedness between tools, thereby compensating for potential information gaps arising from using tool-specific agents.
@@
To get a deeper understanding of the capability of AutoAct, we manually compare the quality of trajectories generated by different methods from five aspects. We ask five NLP experts to individually select the optimal trajectories generated by all methods in terms of the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and overall coherence. The final results are determined based on major votes. During the evaluation, it is hidden for the evaluators of the correspondence between the trajectories and the methods. We delete the reflection-related parts from the trajectories generated by AutoAct and randomly shuffle the order of trajectories of each method in each data to minimize the potential bias as much as possible.
The evaluation results are depicted in Figure 6 and we further provide some cases in Figure 5. We can observe a clear advantage for AutoAct over other methods in determining the action type and action parameters. This indicates that decoupling the missions of planning and tool invocation can lead to better performance for both, alleviating the overwhelming pressure on a single agent. A more intuitive comparison can be observed in Figure 5 (a)(b). AutoAct successfully addresses the failure in ReAct by employing a more scientific combination of tools and making more accurate tool invocations. Furthermore, AutoAct tends to consume more planning rounds than other methods. This allows AutoAct to perform better on harder problems. However, this characteristic can be a double-edged sword when it comes to simple problems. A surprising aspect is that AutoAct can validate its inner (thought) answers by continuing more rounds of self-verification (see Figure 5 (c)). Unfortunately, this can also lead to a longer context, gradually deviating AutoAct from the original question (see Figure 5 (d)).
@@ 6 Related Work
The rise of LLMs has positioned them as the most promising key to unlocking the door to Artificial General Intelligence (AGI), providing robust support for the development of LLM-centered AI agents Wang et al. (2023b); Xi et al. (2023); Wang et al. (2023d, e). Related works focus primarily on agent planning Yao et al. (2023); Song et al. (2022); Chen et al. (2023a), external tools harnessing Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023), collective intelligence among multi-agents Liang et al. (2023); Liu et al. (2023); Chen et al. (2023b), human and social property inside agents Zhang et al. (2023a); Park et al. (2023); Mao et al. (2023), etc. However, despite their success, existing methods still face two major troubles. Firstly, most agents heavily rely on prompts for customization, which makes it difficult to precisely tailor the behavior of the agent, resulting in unexpected performance at times. Secondly, each agent is compelled to master all skills, making it challenging for the agent to achieve expertise in every domain. In response, our approach leverages a proper division-of-labor strategy and fine-tuning each sub-agent to equip different agents with distinct duties. These agents collaborate to accomplish tasks orderly and effectively.
@@
Despite the vast interest in LLM-powered agents, the construction of agents through fine-tuning has received limited attention. Most early works concentrate on fine-tuning to optimize the model’s reasoning capabilities Liu et al. (2022); Fu et al. (2023) or tool proficiency Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023). Recently, Chen et al. (2023a) attempt to fine-tune agents with diverse tasks and trajectories for a better planning capability. Zeng et al. (2023) apply a hybrid instruct-tuning strategy that enhances the agent’s abilities while preserving its generalization. However, these methods still require a model to be a generalist. Moreover, the trajectories in the training data are annotations from GPT-3.5/GPT-4 OpenAI (2022, 2023), which incurs significant costs. Our approach enables the Meta-Agent to autonomously synthesize trajectories and achieve self-planning in a zero-shot manner, without relying on closed-source models or human labor.
7 Conclusion
In this paper, we propose AutoAct, an automatic agent learning framework that does not rely on large-scale annotated data and synthetic trajectories from closed-source models, while alleviating the pressure on individual agents by explicitly dividing the workload, thereby enhancing the collective performance of multi-agents. Experimental results demonstrate that AutoAct performs superior on challenging question-answering benchmarks compared to various strong baselines, and AutoAct with Llama-2-13b model can even obtain comparable performance of zero-shot GPT-3.5-Turbo agent. Extensive analysis indicates the effectiveness of our appropriate division-of-labor strategy, with the trajectory quality generated by AutoAct significantly outperforming that of other methods from multiple aspects.
@@ 7 Conclusion
Tasks and Backbones.
For experimental convenience, we only evaluate our method on complex question-answering tasks with the Llama-2-chat model series. However, there are many other interactive scenarios and backbone models beyond these. Other complex tasks include web Yao et al. (2022); Zhou et al. (2023b), household Puig et al. (2018); Shridhar et al. (2021), robotics Ichter et al. (2022), etc., and more backbone models include Vicuna Zheng et al. (2023), ChatGLM Du et al. (2022), Mistrial Jiang et al. (2023), etc. We plan to conduct further research on applying AutoAct to a wider range of tasks and backbones in the future.
@@
Appendix A Hyper-Parameters
Name Llama-2-7b&13b-chat Llama-2-70b-chat
@@ Appendix A Hyper-Parameters
See Table 4.
Appendix B Task Information
@@
Appendix C Tool Library
See Table 5.

### v4

@@ 4 Results
Compare to Fine-tuning-based Agent Learning Baselines.
Further focusing on FireAct in Tab. 1, despite the aid of GPT-4, FireAct’s approach of assigning the entire planning task to a single model proves to be burdensome. As a result, its performance on ScienceQA even falls short compared to the prompt-based global planning method, Chameleon. AutoAct decouples the planning process and reaches a clear division-of-labor among sub-agents for group planning, resulting in an improvement than FireAct, with $\uparrow$5.77on HotpotQA and $\uparrow$6.67on ScienceQA with Llama-70B model. Additionally, AutoAct achieves self-planning without relying on closed-source models and large-scale labeled datasets, which paves the way for automatic agent learning with open-source models from scratch. In ablation study (§4) and human evaluation (§5), we will further validate that the quality of trajectories synthesized by AutoAct is not inferior to FireAct trained on trajectories synthesized using GPT-4.
Figure 3: Performance of AutoAct on HotpotQA with different training data scales. The {7,13,70}B represents Llama-2-{7,13,70}B-chat models respectively. (a-c) shows the results of the model trained on self-synthesized trajectories. (d-f) represents the results of the model trained on trajectories synthesized by a stronger model, where the dashed line is the baseline trained on self-synthesized trajectories.
Figure 4: Performance of AutoAct on HotpotQA based on different degrees of labor division. One is training a single model with all the differentiated data. Three represents the differentiation into three agents: plan, tool, and reflect. Tool Specified indicates further differentiating the tool-agent with one tool, one agent.
HotpotQA ScienceQA
  AutoAct 48.47 78.61
- reflection 45.66↓2.81 75.28↓3.33
- multi 42.81↓5.66 69.72↓8.89
- fine-tuning 32.84↓15.63 61.94↓16.67
- filtering 32.51↓15.96 59.17↓19.44
@@
Under identical settings, multi-agent architectures generally exhibit better performance than single-agent (ReAct vs. BOLAA, FireAct vs. AutoAct), which aligns with Simon’s theory of bounded rationality. Seemingly contrary to expectations, despite being a single-agent architecture, Chameleon outperforms BOLAA (even FireAct on ScienceQA). However, we analyze that this can be attributed to the way it leverages tools. In Chameleon, the process of deciding tool parameters is considered a form of tool invocation, and specialized few-shot prompts are designed to guide the model through this process. From this aspect, Chameleon, despite nominally a single-agent architecture, exhibits features resembling a multi-agent one, which does not contradict our initial conclusion. Indeed, we can also explain from the perspective of optimizing objectives. Another well-known principle, Goodhart’s Law (Goodhart, 1984), states that “When a measure becomes a target, it ceases to be a good measure”. This implies that optimizing one objective on the same agent will inevitably harm other optimization objectives to some extent. Therefore, it is not optimal to optimize all objectives on a single agent, and a multi-agent architecture happens to address this issue. However, we analyze in §5 that excessive fine-grained division-of-labor is not the best approach.
Figure 5: Case study on HotpotQA. AutoAct (b) successfully addresses the failure in ReAct (a) by employing a more scientific combination of tools and making more accurate tool invocations. With more planning rounds, AutoAct (c) can validate its inner answers by continuing more rounds of self-verification. While this can also lead to a longer context, gradually deviating AutoAct (d) from the original question.
Figure 6: Human evaluation of trajectories generated by Llama-2-70B-chat on HotpotQA. We compare the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and the overall coherence of each trajectory. The figure above displays the Win Rate of each method in each aspect.
@@
Tab. 2 presents the performance of AutoAct on the Llama-70B model after removing certain key processes. It can be observed that the least impactful removal is the - reflect. We investigate that in the zero-shot scenario, the model tends to be over-confident in its answers (as also confirmed in Huang et al. (2024a)). It typically only recognizes its errors when there are obvious formatting mistakes or significant repetitions in the planning process. Consistent with previous findings, the removal of the - multi agents leads to a noticeable decrease in performance. A more exciting discovery is that the results of - multi are comparable to those of FireAct. This indirectly suggests that the trajectory quality generated by the 70B model may be no worse than that of GPT-4. As expected, the performance deteriorates after - fine-tuning, which once again confirms the previous conclusion. To demonstrate the necessity of filtering out planning error data, we specifically remove the filtering process (- filtering) to examine the performance of AutoAct. The results indicate that the damage caused by training on unfiltered data is even greater than that of - fine-tuning.
@@ 5 Analysis
We evaluate the influence of different training data scales on the performance of self-planning with Llama-{7,13,70}B models on HotpotQA in Fig. 3 (a-c). It can be observed that the overall performance of different models goes to stability with minimal waves once the data scale exceeds 200. We speculate that this may be due to the limited ability of naive self-instruct to boost internal knowledge of the language model. As the training data increases, the knowledge which can be extracted through self-instruct decreases. Despite our efforts to filter out duplicate data, the mindless increase can inevitably lead to a significant surge in similar data, which undermines the benefits of increasing the data scale and makes it challenging to improve model performance or even leads to over-fitting. To further confirm the role of training data, we decouple the models from the training data and evaluate their training results on trajectories synthesized by stronger models. From Fig. 3 (d-f), we can see consistent conclusions with previous findings. Therefore, maximizing the diversity of the synthesized data in the database may be a key improvement direction for AutoAct and we leave this for our future work. We can also observe from Fig. 3 (d-e) that the larger the model, the higher the quality of the synthesized data, as the performance of the 7B model shows a gradual increase on self, 13B, and 70B synthesized data.
Moderate division-of-labor benefits group planning performance.
@@
To explore the impact of different granularity of self-differentiation, we further subdivide the tool agent, assigning dedicated agents to manipulate each specific tool. We compare the performance of One agent, Three agents (AutoAct), and the Tool-Specified setting on HotpotQA in Fig. 4. It can be observed that excessive differentiation (Tool-Specified) not only fails to achieve better results but can sometimes even be less effective than not differentiating (One) at all. This is consistent with the findings in Qiao et al. (2023a) which indicate that multi-tool joint learning often outperforms single-tool individual learning. Moreover, it appears that the performance loss of tool-specific agents compared to AutoAct is more significant on harder problems. This is because challenging problems typically require more planning steps and higher levels of collaboration among tools. By unifying tool invocations under one agent, it becomes possible to effectively learn the interconnectedness between tools, thereby compensating for potential information gaps arising from using tool-specific agents. Note the difference from Li et al. (2024), here we are discussing the granularity of division-of-labor among agents with different responsibilities, rather than the voting quantity among mutually equal agents.
@@
To get a deeper understanding of the quality of trajectories generated by different methods, we manually compare them from the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and overall coherence. The detailed human evaluation process can be found in Appx. C. The evaluation results are depicted in Fig. 5&6. We can observe a clear advantage for AutoAct over other methods in the action type and action parameters. This indicates that decoupling the missions of planning and tool invocation can lead to better performance for both, alleviating the overwhelming pressure on a single agent. A more intuitive comparison can be observed in Fig. 5 (a-b). AutoAct successfully addresses the failure in ReAct by employing a more scientific combination of tools and making more accurate tool invocations. Furthermore, AutoAct tends to consume more planning rounds than other methods (the specific average planning rounds is in Appx. D). This allows AutoAct to perform better on harder problems. However, this characteristic can be a double-edged sword when it comes to simple problems. A surprising aspect is that AutoAct can validate its inner answers by continuing more rounds of verification (Fig. 5 (c)). But this can also lead to a longer context, gradually deviating AutoAct from the original question (Fig. 5 (d)).
@@ 6 Related Work
The rise of LLMs has positioned them as the most promising key to unlocking the door to Artificial General Intelligence (AGI), providing robust support for the development of LLM-centered AI agents Wang et al. (2023a); Xi et al. (2023); Wang et al. (2023c, d). Related works focus primarily on agent planning Yao et al. (2023); Song et al. (2022); Chen et al. (2023a), external tools harnessing Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023), collective intelligence among multi-agents Liang et al. (2023); Liu et al. (2023); Chen et al. (2023c), etc. However, despite their success, existing methods still face two major troubles. Firstly, most agents heavily rely on prompts for customization, which makes it difficult to precisely tailor the behavior of the agent, resulting in unexpected performance at times. Secondly, each agent is compelled to master all skills, making it challenging for the agent to achieve expertise in every domain. In response, our approach leverages a proper division-of-labor strategy and fine-tuning each sub-agent to equip different agents with distinct duties. These agents collaborate to accomplish tasks orderly and effectively.
@@
Despite the vast interest in LLM-powered agents, the construction of agents through fine-tuning has received limited attention. Most early works concentrate on fine-tuning to optimize the model’s reasoning capabilities Liu et al. (2022); Fu et al. (2023) or tool proficiency Patil et al. (2023); Qiao et al. (2023a); Qin et al. (2023). Recently, more works have emphasized endowing open-source LLMs with agent capabilities through fine-tuning Chen et al. (2023a); Zeng et al. (2023); Yin et al. (2023); Shen et al. (2024). However, these works suffer from at least one of the following issues: i) the requirement of one single model to be a generalist, ii) the need for a large amount of annotated data, iii) the need for trajectory annotation of closed-source models. Our approach enables the Meta-Agent to synthesize trajectories and achieve a division-of-labor strategy in a zero-shot manner, without relying on closed-source models.
7 Conclusion and Future Work
In this paper, we propose AutoAct, an automatic agent learning framework for QA that does not rely on large-scale annotated data and synthetic trajectories from closed-source models, while alleviating the pressure on individual agents by explicitly dividing the workload. Interesting future directions include: i) expanding AutoAct to more realistic task scenarios (Puig et al., 2018; Zhou et al., 2023a; Xie et al., 2024), ii) boosting more knowledge via self-instruct (as analyzed in §5), iii) iteratively enhancing synthetic trajectories via self-improvement (Huang et al., 2023; Aksitov et al., 2023).
@@ 7 Conclusion and Future Work
Tasks.
In this paper, we mainly focus on complex question-answering tasks. However, there are many other more complex interactive scenarios, including web Yao et al. (2022); Zhou et al. (2023a), household Puig et al. (2018); Shridhar et al. (2021), traveling Xie et al. (2024), robotics Ichter et al. (2022), etc. For example, we have investigated the use of Meta-Agent performing random explorations (Xiang et al., 2023; Murty et al., 2024) in virtual environments to replace the process of task and trajectory synthesis through self-instruct and zero-shot planning. We plan to conduct further research on applying AutoAct to a wider range of tasks based on this in the future.
@@
Ethics Statement
This research was conducted with the highest ethical standards and best practices in research. All our experiments use publicly available datasets (as detailed in §3), avoiding ethical concerns related to privacy, confidentiality, or misuse of personal biological information. The human evaluation process (as detailed in Appx. C) was carried out strictly with fairness and transparency. Consequently, this research is free from any ethical concerns.
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), Tencent AI Lab Rhino-Bird Focused Research Program (RBFR2024003), and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
Appendix A Comparison with Related Works
See Tab. 3
  ReAct Yao et al. (2023) User Prompt Iterative
Reflexion Shinn et al. (2023) User Prompt Iterative
Camel Li et al. (2023) User Prompt Iterative
Chameleon Lu et al. (2023) User Prompt Global
HuggingGPT Shen et al. (2023) User Prompt Global
AutoGPT Torantulino (2023) User Prompt Iterative
BOLAA Liu et al. (2023) User Prompt Iterative
AgentVerse Chen et al. (2023d) User Prompt Iterative
Agents Zhou et al. (2023b) User Prompt Iterative
AgentTuning Zeng et al. (2023) Benchmark GPT-4 Iterative
FireAct Chen et al. (2023a) Benchmark GPT-4 Iterative
Lumos Yin et al. (2023) Benchmark Benchmark + GPT-4 Both
AutoAct (ours) User + Self-Instruct Self-Planning Iterative
Appendix B Baselines and Training Setups
Baselines.
We choose the open-source Llama-2 models Touvron et al. (2023) and Mistral-7B Jiang et al. (2023) as the backbones of our Meta-Agent and sub-agents. The compared baselines are as follows: 1) CoT Wei et al. (2022), the naive Chain-of-Thought reasoning method. 2) ReAct Yao et al. (2023), a well-known single-agent framework based on few-shot learning that performs planning and action iteratively. 3) Chameleon Lu et al. (2023), another few-shot single-agent framework that performs planning before action. 4) Reflexion Shinn et al. (2023), a single-agent framework to reinforce language agents through linguistic feedback. 5) BOLAA Liu et al. (2023), a multi-agent framework that customizes different agents through prompts. 6) ReWOO Xu et al. (2023), a multi-agent framework that decouples reasoning from observations. 7) FireAct Chen et al. (2023a), a single-agent framework with fine-tuning on diverse kinds of trajectories generated by GPT-4 OpenAI (2023). 8) GPT-3.5-Turbo OpenAI (2022). To ensure fairness, we maintain an equal training trajectory volume of 200 for FireAct and AutoAct (200 synthesized data). As Reflexion provides answer correctness labels during reflection but other methods including AutoAct do not, we test all the other methods twice and choose the correct one for evaluation. For all the prompt-based baselines, we uniformly provide two examples in the prompt.
Training Setups.
We fine-tune all our models with LoRA Hu et al. (2022) in the format proposed in Alpaca Taori et al. (2023). Our fine-tuning framework leverages FastChat Zheng et al. (2023) using DeepSpeed Rasley et al. (2020). We detail the hyper-parameters for training in Tab. 4.
Name Mistral-7B&Llama-2-{7,13}B-chat Llama-2-70B-chat
@@ Appendix B Baselines and Training Setups
Appendix C Detailed Process of Human Evaluation
To get a deeper understanding of the capability of AutoAct, we manually compare the quality of trajectories generated by different methods from five aspects. We ask five NLP volunteers to individually select the optimal trajectories generated by all methods in terms of the number of planning rounds, the logical correctness of thoughts, action types, action parameters, and overall coherence. The final results are determined based on major votes. During the evaluation, it is hidden for the evaluators of the correspondence between the trajectories and the methods. We delete the reflection-related parts from the trajectories generated by AutoAct and randomly shuffle the order of trajectories of each method in each data to minimize the potential bias as much as possible.
Appendix D Average Planning Rounds
We compare the planning rounds of AutoAct with various baselines. The win rate of each method is listed in Fig. 6 and comprehensive analysis can be found in §5. Here we present the average planning rounds of various methods on HotpotQA with Llama-2-70B-chat in Tab. 5. Note that to maintain fairness, we exclude the planning steps related to reflection of AutoAct.
Method Easy Medium Hard
  ReAct 3.83 4.02 4.13
BOLAA 3.60 3.76 3.96
FireAct 3.01 3.17 3.70
AutoAct 4.62 4.73 4.96
Table 5: Average planning rounds of various methods on HotpotQA with Llama-2-70B-chat.
Appendix E Task Information
@@
Appendix F Tool Library
To facilitate our agents in automatic task planning, we provide a comprehensive tool library that contains 15 commonly used tools for various complex question-answering tasks. A part of our tools and their corresponding information can be found in Tab. 6. Users can also have the option to expand the tool library according to their specific needs, allowing for more flexible utilization.